# Danny Halawi4, Fred Zhang4, Philip E. Tetlock2,5

### v4

@@ 6 Discussion
We introduced ForecastBench, a dynamic and continuously updated benchmark for evaluating automated forecasting systems. By focusing exclusively on questions that are unresolved at the time of submission, we eliminate the risks of data leakage and ensure a robust evaluation environment. Our initial results demonstrate that while state-of-the-art LLMs exhibit promising potential, they underperform expert human forecasters. This performance gap highlights the challenges in leveraging current LLMs for accurate, real-time forecasting.
We produce a public leaderboard listing the real-time accuracy of top LLMs and humans as well as a standardized dataset of forecasting questions and rationales. Future work should leverage this auxiliary dataset of predictions and rationales to fine-tune models, explore new architectures, and develop adaptive systems better suited for general reasoning in dynamic, real-world environments. Ultimately, ForecastBench serves as a step toward harnessing the full potential of AI-based systems and forecasting in decision-making.
@@ Appendix A Licensing and maintenance
[…]
@@
Our datasets, distributed under the CC BY-SA 4.0 license, will be made available and updated regularly on www.forecastbench.org.
@@
The code underlying our automated system runs on Google Cloud Platform and is available at https://github.com/forecastingresearch/forecastbench under the MIT license.
@@
We ensure the long-term availability and maintenance of the benchmark as it is funded by Open Philanthropy until mid-2027. In case no further funding is provided beyond that point, we will upload our final datasets to GitHub and Hugging Face.
@@ B.1 Question and resolution sets
The latest value of the market or comparison value the day the question was frozen. string
@@
id A unique identifier string given source. A list if this was a combination question. string | array<string>
@@
direction If id has an array value, this is an array of the same length. Each entry $\in\{-1,1\}$. If the value is $1$, it means the question was asked in the normal direction. If the value is $-1$, it means the question was negated in the question. e.g. for a question asking for $P(\neg Q1\cap Q2)$, the value would be $[-1,1]$; all possible directions for this question would be: $[1,1]$, $[-1,1]$, $[1,-1]$, $[-1,-1]$. The value is null when id is a string. string
@@ B.1.1 Examples
id ["meteofrance_TEMPERATURE_celsius.07117.D", "meteofrance_TEMPERATURE_celsius.07240.D"]
@@ B.2.1 General public forecast set
Data is collected from human forecasters via Qualtrics. We collect the reasoning behind forecasts and manually anonymize the data to ensure there is no personally identifiable information in the dataset before posting it online.
@@
id A unique identifier string given source. If instead of a string it’s a list of strings, then this is a combination question. string | array<string>
@@
direction If id has an array value, this is an array of the same length. Each entry is an integer $\in\{-1,1\}$. If the value is $1$, it means the question was asked in the normal direction. If the value is $-1$, it means the question was negated in the question. e.g. for a question asking for $P(\neg Q1\cap Q2)$, the value would be $[-1,1]$. All possible values are: $[1,1]$, $[-1,1]$, $[1,-1]$, $[-1,-1]$, and null. array<number> |null
user_id A randomly generated string associated with the human respondent who submitted the forecast. This value contains no information which could identify said participant and was assigned to the dataset after personal identifiers had been removed. string
suspected_LLM true if we suspect the rationale was created by an LLM, false otherwise. boolean
@@ B.2.2 Superforecaster forecast set
searches An array of search terms used in researching the topic. array<string>
consulted_urls A list of useful urls array<string>
@@ C.2 Categories
Arts & Recreation 0 57 10 16 0 0 0 0 0 83
Economics & Business 1 22 76 35 0 0 166 0 504 804
Environment & Energy 0 3 44 5 0 52 0 0 0 104
Healthcare & Biology 0 8 83 1 0 0 0 214 0 306
Politics & Governance 6 51 208 428 0 0 0 0 0 693
Science & Tech 5 101 200 21 0 0 0 1 0 328
Security & Defense 4 15 157 9 3,150 0 0 0 0 3,335
Sports 0 77 21 4 0 0 0 120 0 222
Other 0 0 1 0 0 0 0 0 0 1
Total 16 334 800 519 3150 52 166 335 504 5,876
@@ D.1 Instructions
Participants in the public survey were first prompted to read a consent form detailing the tasks involved in the study and potential risks to participants, alongside estimates of payment for completing each study (\$5 + bonus payments for forecasting performance in the introductory survey; \$10 for completing the primary survey). In said consent form, the overall task was described as follows: "You will be asked to read some material, to follow some instructions, and to provide forecasts for future events." Then, participants were prompted to give their responses to 20 forecasting questions randomly selected from the 200-question subset of questions provided to the LLMs.
Participants were given a brief description of the task ahead before each question as follows:
You are going to be predicting the probability of the answer to the question below being "Yes" (or "resolving positively").
@@
Participants in the superforecaster survey went through a similar initial experience before being moved into the group forecasting stage described in Section 4. Because of the expertise of this set of forecasters, superforecaster participants were also guaranteed a base payment of \$1,000 for participating in the individual and group stages of the experiment as well as bonus payments for individual-level accuracy.
@@ D.2 Screenshots
The following screenshots (Figure 2, Figure 3 show a few of the questions presented to participants in the public study.
@@
Appendix E Details about LLM "Crowd" Baseline
@@ E.1 Models
To construct a crowd baseline that includes diverse candidates, we evaluate models using the most recent forecasting dataset containing cross-domain questions with true resolutions from Halawi et al. (2024). We assess models from the following organizations: OpenAI, Mistral AI, Qwen, Google, Anthropic, and Meta. Using the same scratchpad prompting method from Halawi et al. (2024), we then select the top three models: GPT-4o, Gemini-1.5.Pro, Claude-3.5-Sonnet. See Table 16 for the results.
@@
The highlights indicate the models we decide to use. Subscript numbers denote 2 standard errors. Above the divider line are the new models we evaluate, while below the divider line are the models evaluated by Halawi et al. (2024) on the same dataset using the same prompt.
@@ E.2 Aggregation Methods
With the forecasts generated by the top 3 models selected in Section E.1, we then compare performance of 5 aggregation methods: Median, Trimmed Mean, Geometric Mean, and Geometric Mean of Log Odds (Satopää et al., 2014). See Table 17 for the results.
@@ E.3 Other Details
LLM Parameters. We set the temperature as 0 and the max output token length as 2000.
@@ Appendix F Prompts
In this section, we present the following prompts: zero-shot (Figure 4), scratchpad (Figure 5), and three prompts (Figure 6, Figure 7, Figure 8) written by superforecasters from the Forecasting Research Institute used to construct the LLM "crowd" baseline. Note that the scratchpad with information retrieval prompt is simply the scratchpad with an additional line "We have retrieved the following information for this question: {retrieved_info}" before the instructions begin. For combination questions, we slightly modify each of the above prompts by including the second question’s information (Figure 9). Finally, we present the prompt used to determine whether a question is marked as valid or invalid in Figure 10.
@@
Figure 4: Zero-shot Prompt from Halawi et al. (2024)
@@
Figure 5: Scratchpad Prompt modified from Halawi et al. (2024)
@@ Appendix G Statistical Details
While our pairwise bootstrapped p-values are precise, our statistical tests do ignore one potentially important source of imprecision. We assume that each question is independent of the other questions. In a world where forecasting accuracy on questions is quite correlated within a topic, or where most events are correlated, this may result in us overstating how confidently we can reject the equivalence of different models. We hope to explore this question in future work, but we are somewhat reassured by the fact that our gathering of forecasting questions from diverse domains and sources makes it unlikely that correlated questions would change our interpretation of these results in any meaningful way.
@@ Appendix H Leaderboards: Top 50
We show the best 50 performers on several leaderboards. The first is the leaderboard for the human question set of 200 standard (non-combination) questions:Table 18. Table 19 shows the leaderboard for the full LLM question set of 1,000 questions. Finally, in Table 20, we present the human leaderboard with combination questions included where humans provided forecasts on both components of the combination question. We derive human forecasts for these combination questions by treating each component of the question as independent.
@@
Superforecaster median forecast ForecastBench 0.126 0.059 0.092 [0.073, 0.112] 0%
Claude-3-5-Sonnet-20240620 Anthropic Freeze values Scratchpad 0.139 0.089 0.114 [0.092, 0.136] <0.01 31%
Public median forecast ForecastBench 0.156 0.072 0.114 [0.095, 0.133] <0.001 24%
@@
Claude-3-5-Sonnet-20240620 Anthropic News with freeze values Scratchpad 0.144 0.094 0.119 [0.098, 0.141] <0.001 30%
GPT-4 OpenAI Freeze values Zero shot 0.164 0.079 0.121 [0.100, 0.143] <0.001 32%
Claude-3-5-Sonnet-20240620 Anthropic Freeze values Zero shot 0.145 0.099 0.122 [0.095, 0.149] <0.001 32%
GPT-4-Turbo-2024-04-09 OpenAI Freeze values Zero shot 0.170 0.077 0.124 [0.102, 0.145] <0.001 33%
GPT-4o OpenAI News with freeze values Scratchpad 0.168 0.082 0.125 [0.107, 0.144] <0.001 26%
Claude-3-5-Sonnet-20240620 Anthropic Scratchpad 0.139 0.113 0.126 [0.104, 0.148] <0.001 29%
GPT-4o OpenAI Freeze values Scratchpad 0.163 0.093 0.128 [0.108, 0.148] <0.001 27%
Claude-3-5-Sonnet-20240620 Anthropic News Scratchpad 0.144 0.116 0.130 [0.109, 0.151] <0.001 26%
Claude-3-5-Sonnet-20240620 Anthropic News Superforecaster 2 0.159 0.103 0.131 [0.112, 0.150] <0.001 25%
Mistral-Large-Latest Mistral AI Freeze values Zero shot 0.170 0.096 0.133 [0.110, 0.156] <0.001 24%
Claude-3-5-Sonnet-20240620 Anthropic News Superforecaster 1 0.153 0.113 0.133 [0.112, 0.154] <0.001 26%
GPT-4o OpenAI Scratchpad 0.163 0.103 0.133 [0.116, 0.150] <0.001 25%
Claude-3-Opus-20240229 Anthropic Freeze values Scratchpad 0.163 0.105 0.134 [0.116, 0.152] <0.001 24%
Claude-3-Opus-20240229 Anthropic Freeze values Zero shot 0.173 0.095 0.134 [0.110, 0.158] <0.001 26%
Mistral-Large-Latest Mistral AI Freeze values Scratchpad 0.161 0.109 0.135 [0.118, 0.152] <0.001 23%
GPT-4o OpenAI News Scratchpad 0.168 0.107 0.137 [0.117, 0.158] <0.001 22%
GPT-4-Turbo-2024-04-09 OpenAI Freeze values Scratchpad 0.179 0.104 0.141 [0.119, 0.163] <0.001 27%
Claude-3-5-Sonnet-20240620 Anthropic Zero shot 0.145 0.138 0.141 [0.115, 0.168] <0.001 24%
Llama-3-70b-Chat-Hf Meta Freeze values Scratchpad 0.189 0.095 0.142 [0.126, 0.158] <0.001 25%
Gemini-1.5-Pro Google Scratchpad 0.168 0.116 0.142 [0.124, 0.160] <0.001 23%
Gemini-1.5-Pro Google Freeze values Scratchpad 0.168 0.116 0.142 [0.124, 0.160] <0.001 23%
Gemini-1.5-Pro Google News with freeze values Scratchpad 0.172 0.113 0.142 [0.123, 0.161] <0.001 22%
GPT-4 OpenAI Freeze values Scratchpad 0.178 0.107 0.143 [0.124, 0.162] <0.001 21%
Llama-3-70b-Chat-Hf Meta Freeze values Zero shot 0.181 0.105 0.143 [0.118, 0.168] <0.001 26%
Claude-3-Opus-20240229 Anthropic Scratchpad 0.163 0.124 0.144 [0.126, 0.161] <0.001 22%
GPT-4-Turbo-2024-04-09 OpenAI Zero shot 0.170 0.118 0.144 [0.125, 0.163] <0.001 22%
Claude-3-Opus-20240229 Anthropic News Superforecaster 1 0.161 0.128 0.144 [0.124, 0.165] <0.001 23%
Claude-2.1 Anthropic Freeze values Scratchpad 0.213 0.077 0.145 [0.128, 0.162] <0.001 26%
GPT-4-Turbo-2024-04-09 OpenAI Scratchpad 0.179 0.111 0.145 [0.129, 0.161] <0.001 23%
Gemini-1.5-Flash Google Freeze values Zero shot 0.183 0.107 0.145 [0.118, 0.172] <0.001 24%
Gemini-1.5-Pro Google News Scratchpad 0.172 0.119 0.145 [0.127, 0.164] <0.001 22%
GPT-4o OpenAI Freeze values Zero shot 0.201 0.091 0.146 [0.121, 0.170] <0.001 28%
Mixtral-8x22B-Instruct-V0.1 Mistral AI Freeze values Scratchpad 0.183 0.108 0.146 [0.129, 0.163] <0.001 21%
Gemini-1.5-Pro Google Freeze values Zero shot 0.186 0.106 0.146 [0.120, 0.172] <0.001 25%
Qwen1.5-110B-Chat Qwen Freeze values Zero shot 0.195 0.097 0.146 [0.125, 0.167] <0.001 22%
GPT-4 OpenAI Scratchpad 0.178 0.116 0.147 [0.132, 0.162] <0.001 18%
GPT-4-Turbo-2024-04-09 OpenAI News with freeze values Scratchpad 0.188 0.108 0.148 [0.126, 0.170] <0.001 26%
Mixtral-8x22B-Instruct-V0.1 Mistral AI Freeze values Zero shot 0.189 0.111 0.150 [0.124, 0.176] <0.001 26%
Mistral-Large-Latest Mistral AI Scratchpad 0.161 0.143 0.152 [0.134, 0.169] <0.001 22%
GPT-4 OpenAI Zero shot 0.164 0.141 0.152 [0.134, 0.170] <0.001 24%
Claude-2.1 Anthropic Scratchpad 0.213 0.092 0.152 [0.134, 0.171] <0.001 23%
Imputed Forecaster ForecastBench 0.250 0.056 0.153 [0.136, 0.170] <0.001 28%
Qwen1.5-110B-Chat Qwen Freeze values Scratchpad 0.193 0.116 0.154 [0.137, 0.172] <0.001 20%
Claude-2.1 Anthropic Freeze values Zero shot 0.220 0.090 0.155 [0.133, 0.177] <0.001 27%
Qwen1.5-110B-Chat Qwen News with freeze values Scratchpad 0.185 0.126 0.155 [0.137, 0.174] <0.001 22%
Mixtral-8x22B-Instruct-V0.1 Mistral AI News with freeze values Scratchpad 0.197 0.115 0.156 [0.138, 0.174] <0.001 23%
GPT-4-Turbo-2024-04-09 OpenAI News Scratchpad 0.188 0.125 0.156 [0.136, 0.177] <0.001 24%
@@
1. For resolved questions, predictions are compared against ground truth, while for unresolved questions, predictions are compared to community aggregates.
@@
3. Pct. more accurate than No. 1: The percent of questions where this forecaster had a better overall score than the best forecaster (with rank 1).
4. Pairwise p-value comparing to No. 1 (bootstrapped): The p-value calculated by bootstrapping the differences in overall score between each model and the best forecaster (the group with rank 1) under the null hypothesis that there’s no difference.
@@
Claude-3-5-Sonnet-20240620 Anthropic Freeze values Scratchpad 0.171 0.062 0.116 [0.111, 0.122] 0%
GPT-4-Turbo-2024-04-09 OpenAI Freeze values Scratchpad 0.171 0.065 0.118 [0.112, 0.123] 0.243 42%
GPT-4o OpenAI Freeze values Scratchpad 0.191 0.054 0.123 [0.117, 0.128] <0.001 43%
Gemini-1.5-Pro Google Freeze values Scratchpad 0.162 0.089 0.125 [0.120, 0.131] <0.001 35%
GPT-4o OpenAI News with freeze values Scratchpad 0.192 0.066 0.129 [0.123, 0.134] <0.001 39%
Gemini-1.5-Pro Google News with freeze values Scratchpad 0.167 0.094 0.131 [0.125, 0.136] <0.001 34%
Claude-3-Opus-20240229 Anthropic Freeze values Zero shot 0.190 0.076 0.133 [0.126, 0.139] <0.001 41%
Qwen1.5-110B-Chat Qwen Freeze values Scratchpad 0.176 0.092 0.134 [0.128, 0.139] <0.001 31%
GPT-4-Turbo-2024-04-09 OpenAI Scratchpad 0.171 0.097 0.134 [0.129, 0.139] <0.001 31%

### v5

@@ 6 Discussion
We introduced ForecastBench, a dynamic and continuously updated benchmark for evaluating LLM forecasting capabilities. By focusing exclusively on questions that are unresolved at the time of submission, we eliminate the risks of data leakage and ensure a robust evaluation environment. Our initial results demonstrate that while state-of-the-art LLMs exhibit promising potential, they underperform superforecasters. This performance gap highlights the challenges in leveraging current LLMs for accurate, real-time forecasting.
We produce a public leaderboard listing the real-time accuracy of top LLMs and humans as well as a standardized dataset of forecasting questions and rationales. Future work should leverage this auxiliary dataset of predictions and rationales to fine-tune models, explore new architectures, and develop adaptive systems better suited for general reasoning in dynamic, real-world environments. Ultimately, ForecastBench serves as a step toward harnessing the full potential of AI-based systems for forecasting and decision-making.
7 Reproducibility statement
One reason we’ve open-sourced our code (link in Appendix A) is to allow for independent verification of our results. See Appendix I for reproducing the human forecast sets, Appendix J for reproducing LLM forecast sets, and Appendix K for resolving the forecasts and creating the leaderboard.
8 Ethics statement
Human survey subjects in both the public and superforecaster surveys are made aware prior to their participation in the study via an informed consent form (approved by our IRB, number 855431) that their forecast/rationale data may be publicly released and used to train large language models or other AI systems, with said data carefully reviewed and anonymized.
We have manually reviewed text provided by human participants to ensure that no personally identifiable information is released as part of our human forecast datasets, per IRB requirements. Similar manual reviews of text data will take place as part of every future human forecasting round.
Acknowledgments
We thank Open Philanthropy for providing a grant to fund this work, allowing us to maintain the benchmark at least until mid-2027.
We thank Otto Kuusela for helpful discussions, Sam Glover and Molly Hickman for writing superforecaster prompts, and Harrison Durland and Victoria Schmidt for their research assistance.
@@ Appendix A Licensing and maintenance
The latest leaderboards, updated nightly, are available on www.forecastbench.org. Documentation is available on the repository wiki: github.com/forecastingresearch/forecastbench/wiki.
@@
Our datasets, distributed under the CC BY-SA 4.0 license, are available on www.forecastbench.org/datasets.html. Historical updates to the resolution datasets and leaderboards are available on github.com/forecastingresearch/forecastbench-datasets. Bi-weekly question sets are also released via this repository.
@@
The code underlying our automated system runs on Google Cloud Platform and is available at github.com/forecastingresearch/forecastbench under the MIT license.
Participating
Bi-weekly forecasting rounds are open to LLM teams. Instructions for participating are can be found on the wiki.
@@
We ensure the long-term availability and maintenance of the benchmark as it is funded by Open Philanthropy until mid-2027. If no further funding is provided beyond that point, datasets will continue to be made available on GitHub.
@@ B.1 Question and resolution sets
The latest value of the market or comparison value the day the question set was generated. string
@@
id A unique identifier string given source. If instead of a string it’s an array of strings, then this is a combination question. string | array<string>
@@
direction If id has an array value, this is an array of the same length. Each entry $\in\{-1,1\}$. If the value is $1$, the question was asked in the normal direction. If the value is $-1$, the question was negated in the combination question e.g., for a question asking for $P(\neg Q1\cap Q2)$, the value would be $[-1,1]$; all possible directions for this question would be: $[1,1]$, $[-1,1]$, $[1,-1]$, $[-1,-1]$. The value is null when id is a string. string
@@ B.1.1 Examples
id [“meteofrance_TEMPERATURE_celsius.07117.D”, “meteofrance_TEMPERATURE_celsius.07240.D”]
@@ B.2.1 General public forecast set
Data is collected from human forecasters via Qualtrics. We collect the rationale behind forecasts and manually anonymize the data to ensure there is no personally identifiable information in the dataset before posting it online.
@@
id A unique identifier string given source. If instead of a string it’s an array of strings, then this is a combination question. string | array<string>
@@
direction If id has an array value, this is an array of the same length. Each entry is an integer $\in\{-1,1\}$. If the value is $1$, the question was asked in the normal direction. If the value is $-1$, the question was negated in the combination question e.g., for a question asking for $P(\neg Q1\cap Q2)$, the value would be $[-1,1]$. All possible values are: $[1,1]$, $[-1,1]$, $[1,-1]$, $[-1,-1]$, and null. array<number> |null
user_id A randomly generated string associated with the human respondent who submitted the forecast. This value contains no information which could identify said participant and was assigned to the dataset after personal identifiers had been removed. Only required for participants from the general public and superforecaster surveys. string
@@ B.2.2 Superforecaster forecast set
searches An array of search terms used in researching the topic. array<string> |null
consulted_urls A list of useful URLs array<string> |null
@@ C.2 Categories
Arts & Recreation 0 42 10 65 0 0 0 0 0 117
Economics & Business 2 13 55 154 0 0 166 0 509 899
Environment & Energy 0 2 37 7 0 52 0 0 0 98
Healthcare & Biology 0 8 71 3 0 0 0 215 0 297
Politics & Governance 3 16 128 188 0 0 0 0 0 335
Science & Tech 5 66 172 15 0 0 0 1 0 259
Security & Defense 3 9 109 12 3,220 0 0 0 0 3,353
Sports 0 65 18 468 0 0 0 137 0 688
Other 6 151 122 2 0 0 0 75 0 356
Total 19 372 722 914 3,220 52 166 428 509 6,402
@@ D.1 Instructions
Participants in the public survey were first prompted to read a consent form detailing the tasks involved in the study and potential risks to participants, alongside estimates of payment for completing each study (\$5 + bonus payments for forecasting performance in the introductory survey; \$10 for completing the primary survey). In said consent form, the overall task was described as follows: “You will be asked to read some material, to follow some instructions, and to provide forecasts for future events.” Then, participants were prompted to give their responses to 20 forecasting questions randomly selected from the 200-question subset of questions provided to the LLMs.
Participants were given a brief description of the task before each question:
You are going to be predicting the probability of the answer to the question below being “Yes” (or “resolving positively”).
@@
Participants in the superforecaster survey went through a similar initial experience before being moved into the group forecasting stage described in Section 4. Because of the expertise of this set of forecasters, superforecasters were also guaranteed a base payment of \$1,000 for participating in the individual and group stages of the experiment as well as bonus payments for individual-level accuracy.
@@ D.2 Screenshots
Figure 2 and Figure 3 show a few of the questions presented to participants in the public study.
@@
Appendix E LLM “ensemble” baseline
@@ E.1 Models
To construct an ensemble baseline that includes diverse candidates, we evaluate models using the most recent forecasting dataset containing cross-domain questions with true resolutions from Halawi et al. [2024]. We assess models from the following organizations: OpenAI, Mistral AI, Qwen, Google, Anthropic, and Meta. Using the same scratchpad prompting method from Halawi et al. [2024], we then select the top three models: GPT-4o, Gemini-1.5.Pro, Claude-3.5-Sonnet. See Table 16 for the results.
@@
The highlights indicate the models we decide to use. Subscript numbers denote 2 standard errors. Above the divider line are the new models we evaluate, while below the divider line are the models evaluated by Halawi et al. [2024] on the same dataset using the same prompt.
@@ E.2 Aggregation methods
With the forecasts generated by the top 3 models selected in Section E.1, we then compare performance of 5 aggregation methods: Median, Trimmed Mean, Geometric Mean, and Geometric Mean of Log Odds [Satopää et al., 2014]. See Table 17 for the results.
@@ E.3 Other details
LLM Parameters. We set the temperature to 0 and the max output token length to 2000.
@@ Appendix F Prompts
In this section, we present the following prompts: zero-shot (Figure 4), scratchpad (Figure 5), and three prompts (Figure 6, Figure 7, Figure 8) written by superforecasters from the Forecasting Research Institute used to construct the LLM “ensemble” baseline. Note that the scratchpad with information retrieval prompt is simply the scratchpad with an additional line “We have retrieved the following information for this question: {retrieved_info}” before the instructions begin. For combination questions, we slightly modify each of the above prompts by including the second question’s information (Figure 9). Finally, we present the prompt used to determine whether a question is marked as valid or invalid in Figure 10.
@@
Figure 4: Zero-shot Prompt from Halawi et al. [2024]
@@
Figure 5: Scratchpad Prompt modified from Halawi et al. [2024]
@@ Appendix G Statistical details
While our pairwise bootstrapped $p$-values are precise, our statistical tests ignore one potentially important source of imprecision. We assume that each question is independent of the other questions. In a world where forecasting accuracy is quite correlated within a topic, or where most events are correlated, this may result in us overstating how confidently we can reject the equivalence of different models. We hope to explore this question in future work, but we are somewhat reassured by the fact that our gathering of forecasting questions from diverse domains and sources makes it unlikely that correlated questions would change our interpretation of these results in any meaningful way.
@@ Appendix H Leaderboards: top 50
We show the best 50 performers on several leaderboards. In Table 18we show the leaderboard for the human question set of 200 standard (non-combination) questions. Table 19 shows the leaderboard for the full LLM question set of 1,000 questions. Finally, in Table 20, we present the human leaderboard with combination questions included where humans provided forecasts on both components of the combination question. We derive human forecasts for these combination questions by treating each component of the combination question as independent.
@@
Superforecaster median forecast ForecastBench 0.118 0.074 0.096 [0.076, 0.116] 0%
Public median forecast ForecastBench 0.153 0.089 0.121 [0.101, 0.141] <0.001 22%
Claude-3-5-Sonnet-20240620 Anthropic Freeze values Scratchpad 0.138 0.107 0.122 [0.099, 0.146] <0.001 31%
@@
Claude-3-5-Sonnet-20240620 Anthropic News with freeze values Scratchpad 0.142 0.112 0.127 [0.104, 0.150] <0.001 29%
GPT-4-Turbo-2024-04-09 OpenAI Freeze values Zero shot 0.162 0.095 0.128 [0.105, 0.151] <0.001 32%
Claude-3-5-Sonnet-20240620 Anthropic Freeze values Zero shot 0.145 0.117 0.131 [0.103, 0.159] <0.001 31%
GPT-4 OpenAI Freeze values Zero shot 0.167 0.096 0.132 [0.109, 0.155] <0.001 31%
GPT-4o OpenAI News with freeze values Scratchpad 0.162 0.105 0.133 [0.113, 0.154] <0.001 25%
Claude-3-5-Sonnet-20240620 Anthropic Scratchpad 0.138 0.133 0.136 [0.113, 0.158] <0.001 28%
GPT-4o OpenAI Freeze values Scratchpad 0.161 0.113 0.137 [0.115, 0.158] <0.001 27%
Claude-3-5-Sonnet-20240620 Anthropic News Scratchpad 0.142 0.137 0.139 [0.117, 0.161] <0.001 26%
Claude-3-Opus-20240229 Anthropic Freeze values Zero shot 0.163 0.115 0.139 [0.114, 0.164] <0.001 23%
Claude-3-5-Sonnet-20240620 Anthropic News Superforecaster 2 0.158 0.123 0.140 [0.120, 0.161] <0.001 25%
GPT-4o OpenAI Scratchpad 0.161 0.125 0.143 [0.125, 0.160] <0.001 24%
Claude-3-5-Sonnet-20240620 Anthropic News Superforecaster 1 0.150 0.135 0.143 [0.120, 0.166] <0.001 25%
GPT-4o OpenAI News Scratchpad 0.162 0.126 0.144 [0.122, 0.165] <0.001 22%
Claude-3-Opus-20240229 Anthropic Freeze values Scratchpad 0.160 0.129 0.144 [0.125, 0.163] <0.001 23%
Mistral-Large-Latest Mistral AI Freeze values Zero shot 0.173 0.117 0.145 [0.121, 0.169] <0.001 23%
Gemini-1.5-Pro Google News with freeze values Scratchpad 0.163 0.130 0.146 [0.127, 0.165] <0.001 23%
Gemini-1.5-Pro Google Scratchpad 0.162 0.131 0.147 [0.129, 0.164] <0.001 23%
Mistral-Large-Latest Mistral AI Freeze values Scratchpad 0.161 0.133 0.147 [0.129, 0.165] <0.001 22%
GPT-4-Turbo-2024-04-09 OpenAI Zero shot 0.162 0.137 0.149 [0.129, 0.170] <0.001 23%
GPT-4-Turbo-2024-04-09 OpenAI Freeze values Scratchpad 0.176 0.123 0.150 [0.126, 0.173] <0.001 26%
GPT-4 OpenAI Freeze values Scratchpad 0.174 0.126 0.150 [0.130, 0.170] <0.001 21%
Gemini-1.5-Pro Google Freeze values Scratchpad 0.162 0.138 0.150 [0.132, 0.169] <0.001 22%
Claude-3-5-Sonnet-20240620 Anthropic Zero shot 0.145 0.155 0.150 [0.123, 0.177] <0.001 24%
Gemini-1.5-Pro Google News Scratchpad 0.163 0.141 0.152 [0.133, 0.171] <0.001 22%
Claude-3-Opus-20240229 Anthropic News Superforecaster 1 0.157 0.147 0.152 [0.131, 0.174] <0.001 22%
GPT-4-Turbo-2024-04-09 OpenAI News with freeze values Scratchpad 0.178 0.128 0.153 [0.130, 0.175] <0.001 25%
GPT-4o OpenAI Freeze values Zero shot 0.197 0.109 0.153 [0.128, 0.178] <0.001 26%
Llama-3-70b-Chat-Hf Meta Freeze values Scratchpad 0.191 0.116 0.153 [0.135, 0.171] <0.001 25%
Mixtral-8x22B-Instruct-V0.1 Mistral AI Freeze values Scratchpad 0.181 0.127 0.154 [0.137, 0.172] <0.001 21%
GPT-4-Turbo-2024-04-09 OpenAI Scratchpad 0.176 0.133 0.154 [0.137, 0.171] <0.001 23%
Gemini-1.5-Pro Google Freeze values Zero shot 0.185 0.124 0.155 [0.127, 0.182] <0.001 25%
Claude-3-Opus-20240229 Anthropic Scratchpad 0.160 0.149 0.155 [0.136, 0.173] <0.001 22%
GPT-4 OpenAI Scratchpad 0.174 0.137 0.155 [0.140, 0.171] <0.001 18%
Qwen1.5-110B-Chat Qwen Freeze values Zero shot 0.197 0.115 0.156 [0.134, 0.179] <0.001 20%
Claude-2.1 Anthropic Freeze values Scratchpad 0.217 0.095 0.156 [0.138, 0.175] <0.001 24%
Gemini-1.5-Flash Google Freeze values Zero shot 0.191 0.125 0.158 [0.130, 0.186] <0.001 23%
Mixtral-8x22B-Instruct-V0.1 Mistral AI Freeze values Zero shot 0.185 0.131 0.158 [0.131, 0.185] <0.001 24%
Llama-3-70b-Chat-Hf Meta Freeze values Zero shot 0.194 0.124 0.159 [0.133, 0.185] <0.001 25%
GPT-4-Turbo-2024-04-09 OpenAI News Scratchpad 0.178 0.144 0.161 [0.140, 0.182] <0.001 24%
Imputed Forecaster ForecastBench 0.250 0.073 0.162 [0.142, 0.181] <0.001 26%
Qwen1.5-110B-Chat Qwen Freeze values Scratchpad 0.183 0.141 0.162 [0.144, 0.180] <0.001 18%
Claude-2.1 Anthropic Scratchpad 0.217 0.109 0.163 [0.143, 0.183] <0.001 22%
Qwen1.5-110B-Chat Qwen News with freeze values Scratchpad 0.179 0.148 0.163 [0.144, 0.183] <0.001 22%
Claude-2.1 Anthropic Freeze values Zero shot 0.221 0.108 0.164 [0.141, 0.188] <0.001 27%
Mistral-Large-Latest Mistral AI Scratchpad 0.161 0.168 0.165 [0.146, 0.183] <0.001 22%
Claude-3-5-Sonnet-20240620 Anthropic News Scratchpad 0.197 0.133 0.165 [0.144, 0.186] <0.001 19%
Claude-3-Opus-20240229 Anthropic Zero shot 0.163 0.167 0.165 [0.140, 0.191] <0.001 18%
@@
1. Shows performance on the 200 standard questions provided in the human question set at the 7-, 30-, 90-, and 180-day forecast horizons.
2. The full leaderboard is available at www.forecastbench.org. Online results are updated nightly, so may be slightly different than the version presented here.
3. For resolved market questions, forecasts are compared against ground truth while for unresolved market questions, they are compared to community aggregates.
@@
5. Pairwise $p$-value comparing to No. 1 (bootstrapped): The $p$-value calculated by bootstrapping the differences in overall score between each model and the best
forecaster under the null hypothesis that there’s no difference.
6. Pct. more accurate than No. 1: The percent of questions where this forecaster had a better overall score than the best forecaster.
@@
Claude-3-5-Sonnet-20240620 Anthropic Freeze values Scratchpad 0.169 0.078 0.123 [0.117, 0.129] 0%
GPT-4-Turbo-2024-04-09 OpenAI Freeze values Scratchpad 0.172 0.080 0.126 [0.120, 0.132] 0.096 43%
GPT-4o OpenAI Freeze values Scratchpad 0.186 0.069 0.128 [0.122, 0.133] <0.01 43%
Gemini-1.5-Pro Google Freeze values Scratchpad 0.162 0.106 0.134 [0.128, 0.139] <0.001 35%
GPT-4o OpenAI News with freeze values Scratchpad 0.190 0.084 0.137 [0.131, 0.143] <0.001 39%
Gemini-1.5-Pro Google News with freeze values Scratchpad 0.166 0.111 0.139 [0.133, 0.144] <0.001 34%
Claude-3-Opus-20240229 Anthropic Freeze values Zero shot 0.186 0.093 0.139 [0.133, 0.146] <0.001 41%
Qwen1.5-110B-Chat Qwen Freeze values Scratchpad 0.176 0.108 0.142 [0.136, 0.148] <0.001 30%